3x+2y=85
5x+3y=130

Answer by Alan3354(69443)

3x+2y=85
5x+3y=130
-------------
Multiply the 1st eqn by e
3x+2y=85
---
9x + 6y = 255
================
Multiply the 2nd eqn by 2, then subtract.